OpenStudy (anonymous):

2 + (a + b) = (2 + a) + b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

Wait this is the associative property

OpenStudy (unklerhaukus):

Associative Property of Addition (x+y)+z=x+(y+z)

OpenStudy (anonymous):

You're associating diff terms or grouping different terms together that's why it's the assoociative prop

OpenStudy (unklerhaukus):

Distributive Property\[x(y+z)=xy+xz\]